1. Estimated demand function for processed pork is Q=171-20p+20pb+3pc+2y. using the estimated demand function for processed pork , show how the quantity demanded at a given price changes as per capital income, Y, increases by N100? (6 marks)

2.If the supply of corn is Qn= a + bp and the supply by the rest of the word is Qr=c+ep, what is the word supply? what is the world inverse supply? (6 marks)

3. The demand function for rose is Q=a-pb, and the supply function is Q=c+ep+ft, where a, b, c,e, and f are positive constants and t is the average temperature in a month. show how the equilibrium quantity and price vary with temperature? (8 marks)
1
Expert's answer
2015-05-06T09:56:56-0400
1. Q=171-20p+20pb+3pc+2y
If per capital income, Y, increases by N100, the quantity demanded at a given price will increase by 2*100 = 200 units.
2. If the supply of corn is Qn= a + bp and the supply by the rest of the word is¬ Qr=c+ep, the world supply is Qw = Qn + Qr = a + c + (b + e)*p.
3. Qd = a¬pb, Qs = c+ep+ft. In equilibrium Qd = Qs, so:
a¬pb = c+ep+ft
p*(e + b) = a - c - ft
p = (a - c - ft)/(e + b)
So, equilibrium price increases, when the temperature falls and equilibrium quantity falls too.